# Collective Outlier Detection and Enumeration with Conformalized Closed Testing

### v2

@@ 2.3.2 Shirashi’s Locally Most Powerful Rank Test
In this section, let us imagine the calibration scores $X_{1},\ldots,X_{m}$ are drawn independently from an unknown distribution $F$ and the test scores $Y_{1},\ldots,Y_{n}$ are drawn independently from a mixture distribution $(1-\theta)F+\theta G(F)$, where $\theta\in[0,1]$ is the unknown proportion of outliers and $G$ is a distribution function on $[0,1]$. As a starting point, let us assume $G$ is known. In practice, however, the test discussed in this section can be implemented using a data-driven estimate $\hat{G}$ of $G$, as detailed in Appendix A2. Importantly, we will show that this test remains valid when using $\hat{G}$ in place of $G$, provided that $\hat{G}$ is estimated in a manner that preserves the exchangeability between the calibration and test scores.
@@
For any $S\subseteq[n]$ and any $j\in S$, let $R_{S}^{j}$ denote the rank of $Y_{j}$ among $(X_{1},\ldots,X_{m},(Y_{j})_{j\in S})$. For any distribution function $G$ on $[0,1]$ with density function $g$, define the statistic $T^{\text{G}}_{S}$ as:
$T^{\text{G}}_{S}$ $=\sum_{j\in S}\mathbb{E}[g(U_{m+s}^{(R_{S}^{j})})],$ (4)
@@
where $U_{m+s}^{(R_{S}^{j})}$ is the $R_{S}^{j}$-th order statistic in a sample of Uniform random variables of size $m+s$ and $s=|S|$. The test rejects $H_{S}=\cap_{j\in S}H_{j}$ at level $\alpha\in(0,1)$ if $T^{\text{G}}_{S}$ is larger than a suitable critical value $c^{\text{G}}_{\alpha}(m,s)$, discussed later. The indicator of this rejection event is:
$\phi^{\mathrm{G}}_{S}$ $=\mathds{1}\left\{T^{\text{G}}_{S}>c^{\text{G}}_{\alpha}(m,s)\right\}.$ (5)
This reduces to the classic WMW test statistic in the special case of $G(F)=F^{2}$. Lehmann (1953) investigated the power of rank tests in the case when the distribution of the scores assigned to outliers in the test set is equal to $G(F)=F^{k}$ for any integer $k\geq 2$, which is referred to as Lehmann’s alternative. Here, $F^{k}$ represents the distribution of the maximum of $k$ independent random variables with distribution $F$. If $X\sim F$ and $Y\sim F^{k}$, then $\mathbb{P}(X<Y)=k/(k+1)$. It is worth noting that the distance of $F^{k}$ from $F$ remains the same for any specification of $F$. Lehmann (1953) proved that in the mixture model $Y_{i}\sim(1-\theta)F+\theta F^{k}$ with $k=2$, the classical WMW test that rejects for large values of
$T_{S}^{\text{WMW}}=\sum_{j\in S}R_{S}^{j}$
is the locally most powerful invariant (LMPI) test of $H_{0}:\theta=0$, i.e., it is invariant under the group of continuous and strictly increasing transformations of the scores. Shiraishi (1985) extended the result by Lehmann (1953) and proved that the locally most powerful rank test for testing $H_{0}$ given any continuous distribution function $G$ is the one in Equation (5).
@@
[(Shiraishi, 1985)] Assume $X_{1},\ldots,X_{m}$ i.i.d. with continuous c.d.f. $F$, and $Y_{1},\ldots,Y_{n}$ i.i.d. with c.d.f. $(1-\theta)F+\theta G(F)$ for some $\theta\in[0,1]$ and some continuous c.d.f. $G$ on $[0,1]$ with density function $g$. Denote the combined sample $(X_{1},\ldots,X_{m},Y_{1},\ldots,Y_{n})$ as $(W_{1},\ldots,W_{N})$, with order statistics $W_{(1)}<\ldots<W_{(N)}$ and $N=m+n$. Denote as $R_{j}$ the rank of $Y_{j}$ in this sequence. Then, the LMPI test rejects $H_{0}:\theta=0$ for large values of:
$T^{\mathrm{G}}=\sum_{j\in[n]}\mathbb{E}[g(U_{N}^{(R_{j})})].$
Under the Lehmann’s alternative $G(F)=F^{k}$ for integers $k\geq 2$ the associated density function is $g(u)=ku^{k-1}\mathds{1}[u\in[0,1]]$ and the test statistics in Equation (4) reduce to the collection depending on $k$:
$T^{\text{WMW, k}}_{S}$ $=\frac{k}{(N+1)\cdot\ldots\cdot(N+k-1)}\sum_{j\in S}\prod_{l=0}^{k-2}(R_{S}^{j}+l),$ (6)
which are the locally most powerful rank tests against the class of Lehmann’s alternatives.
For $k=2$, the test statistic in Equation (6) differs from the WMW statistic only by a multiplicative constant. It is worth noting that the Mann-Whitney U-statistic can be expressed as $\sum_{j\in[n]}R_{j}-n(n+1)/2$, differing only by a constant from the WMW statistic. As a consequence, inference results provided by these three statistics are equivalent.
The critical value $c^{\text{G}}_{\alpha}(m,s)$ in (5) can be obtained either through a Monte Carlo simulation of the permutation distribution—assuming only the exchangeability of the null scores and feasible for small sample sizes—or via an asymptotic approximation (Shiraishi, 1985) in the limit of large sample sizes, under the additional assumption of independence.
@@
[(Shiraishi, 1985)] In the setting of Theorem 1 the rescaled LMPI test statistic
$\frac{\sqrt{N(N-1)}(T^{\mathrm{G}}-n\mu_{N})}{\sqrt{nm\sum_{h\in[N]}(\mathbb{E}[g(U_{N}^{(h)})]-\mu_{N})^{2}}}$ (7)
with $\mu_{N}=N^{-1}\sum_{h\in[N]}\mathbb{E}[g(U_{N}^{(h)})]$ has asymptotically a standard normal distribution under the null hypothesis $H_{0}$, as $m,n\to\infty$.
As a consequence, when considering any $S\subseteq[n]$ and denoting by $s$ its cardinality, the asymptotic distribution of the test statistic $T^{G}_{S}$ is asymptotically normal with mean
$\mu_{m+s}=\frac{s}{m+s}\sum_{h\in[m+s]}\mathbb{E}[g(U^{(h)}_{m+s})]$
and variance
$\sigma^{2}_{m+s}=\frac{sm}{m+s(m+s-1)}\sum_{h\in[m+s]}\left(\mathbb{E}[g(U^{(h)}_{m+s})]-\frac{\mu_{m+s}}{s}\right)^{2}.$
We conclude this section by emphasizing that the oracle procedure studied in Theorem 2 can be translated into a practical test by replacing the unknown alternative distribution $G$ with a suitable empirical estimate $\hat{G}$, as explained in Appendix A2. As shown in Section 3, in some scenarios of interest this data-driven approximation of the oracle LMPI test can achieve much higher power compared to other local testing methods, such as Simes’, Fisher’s, and the WMW test. In any case, the true strength of our ACODE method lies in its flexibility, as it does not depend on any single testing procedure but can instead dynamically select the most effective approach based on the data at hand.
@@ 2.4 Estimating the Number of Outliers via Closed Testing
Input : Individual hypotheses $H_{j}$, for all $j\in[n]$; a method $\phi$ for carrying out local tests $\phi_{S}$ of $H_{S}$, for any $S\subseteq[n]$; significance level $\alpha\in(0,1)$.
@@
2For each $S\subseteq[n]$, test $H_{S}$ at level $\alpha$; let $\phi_{S}\in\{0,1\}$ denote the rejection indicator.
3For each $S\subseteq[n]$, adjust the local test by setting $\bar{\phi}_{S}:=\min\{\phi_{K}:K\supseteq S\}$.
4For each $S\subseteq[n]$, the lower bound for the number of true discoveries in $S$ is:
@@
Output : A $(1-\alpha)$ lower bound $d(S)$ for the number of outliers in any $S\subseteq[n]$.
@@
The first step of Algorithm 1 tests all possible intersection hypotheses $H_{S}$. We refer to these tests as the local tests. In the second step, a multiplicity adjustment is applied to all local tests. The latter results in $H_{S}$ being rejected, setting $\bar{\phi}_{S}=1$, if and only if $H_{K}$ was locally rejected for all supersets $K$ of $S$. Finally, $d(S)$ is given by (8). This is easiest to explain in the special case of $S=[n]$, when the lower confidence bound for the total number of outliers is given by the difference between the total number of test points, $n$, and the cardinality of the largest subset for which the multiplicity-adjusted local test failed to reject the null hypothesis. For example, if the global null $H_{[n]}$ is rejected but there exists a subset $K\subseteq[n]$ with cardinality $|K|=n-1$ for which $\phi_{K}=0$, then $d([n])=1$.
@@
If Algorithm 1 is applied using a valid local test $\phi$ satisfying $\mathbb{P}[\phi_{I_{0}}=1\mid H_{I_{0}}]\leq\alpha$, then the output lower bounds $d(S)$ satisfy:
@@
The availability of computational shortcuts for Algorithm 1 hinges on the choice of local tests (Goeman et al., 2019, 2021; Tian et al., 2023). See Appendix A3 for a review of relevant shortcuts and implementation details. Appendix A3.1 reviews the shortcut of Goeman et al. (2019), which reduces the cost of Algorithm 1 to $\mathcal{O}(n\log n)$ in the case of local Simes tests. Appendix A3.3 presents a shortcut for an adaptive version of the Simes local test. Appendix A3.4 introduces a shortcut for Shirashi’s LMPI test described in Section 2.3.2. Appendix A3.5 reviews the strategy of Tian et al. (2023), which provides similar shortcuts for a family of local tests including Fisher’s method and the WMW test.
@@ 2.5 Conformalized Closed Testing
Valid conformity scores can be obtained through various methods. A common approach involves using a one-class classifier trained on an independent set of inliers (Bates et al., 2023). Alternatively, one can employ a binary classifier via positive-unlabeled learning (Marandon et al., 2024). Another option arises when labeled outliers are available; here, the scores can be generated using a binary classifier as discussed in Liang et al. (2024). For simplicity, we focus on scenarios where all labeled data consist of inliers.
@@
8 Apply $\mathcal{A}$ to evaluate the calibration and test scores, $(X_{1},\ldots,X_{m})$ and $(Y_{1},\ldots,Y_{n})$.
9
10For any desired $S\subseteq[n]$, compute $d(S)$ using Algorithm 1 (ref. Appendix A3).
@@
Output : A $(1-\alpha)$ lower bound $d(S)$ for the number of outliers in any $S\subseteq[n]$.
@@
Proposition 1 implies Algorithm 2 produces valid inferences as long as all local tests are valid, and our method is precisely designed to achieve this (ref. Section 2.3). While the lower bound $d(S)$ output by Algorithm 2 is simultaneously valid for all possible subsets $S\subseteq[n]$, in the sense of (9), this does not mean Algorithm 2 needs to explicitly output $d(S)$ for all $S\subseteq[n]$. On the contrary, one would typically apply Algorithm 2 focusing on a particular (but possibly data-driven) choice of $S$, as demonstrated in Section 3. In any case, the components of Algorithm 2 involving the model training and the computation of the conformity scores only need to be applied once, irrespective of $S$.

### v3

@@ 2.3.2 Shirashi’s Locally Most Powerful Rank (LMPR) Test
As a starting point, let us assume $g$ is known. In practice, however, the test discussed in this section can be implemented using a data-driven estimate $\hat{g}$ of $g$, as detailed in Appendix A2. Importantly, we will show in Proposition A3 that this test remains valid when using $\hat{g}$ in place of $g$, provided that $\hat{g}$ is estimated in a manner that preserves the exchangeability between the calibration and test scores.
@@
Let $R_{1},\ldots,R_{N}$ denote the ranks of the pooled sample $(X_{1},\ldots,X_{m},Y_{1},\ldots,Y_{n})$, where $N=m+n$. For any distribution $G$ on $[0,1]$ with density function $g$, define the statistic
$T^{\text{g}}$ $=\sum_{j=m+1}^{N}\mathbb{E}[g(U_{N}^{(R_{j})})],$ (4)
@@
where $U_{N}^{(R_{j})}$ is the $R_{j}$-th order statistic in a sample of Uniform random variables of size $N$. The test rejects $H_{0}:\theta=0$ at level $\alpha\in(0,1)$ if $T^{\text{g}}$ is larger than a suitable critical value $c^{\text{g}}_{\alpha}(m,n)$, discussed later. The indicator of this rejection event is:
$\phi^{\mathrm{g}}$ $=\mathds{1}\left\{T^{\text{g}}>c^{\text{g}}_{\alpha}(m,n)\right\}.$ (5)
In the special case of $G(F)=F^{2}$, the statistic $T^{\text{g}}$ in (4) reduces to $\frac{2}{N+1}\cdot T^{\text{WMW}}$, where
$T^{\text{WMW}}=\sum_{j=m+1}^{N}R_{j}$
is the classical WMW statistic. The corresponding test is therefore equivalent to the WMW test, since the multiplicative constant can be ignored. Lehmann (1953) proved that in the mixture model $Y_{i}\sim(1-\theta)F+\theta F^{2}$, the WMW test is the LMPR test of $H_{0}:\theta=0$, i.e. it is most powerful within the class of all rank tests at level $\alpha$ uniformly in a small neighbourhood of 0. Shiraishi (1985) extended the result by Lehmann (1953) and proved that the LMPR test for $H_{0}$ given any continuous density function $g$ is the one in Equation (5); see Theorem A1 in Appendix A1.5.
In Appendix A1.5, we study the power of Shiraishi’s LMPR test and compare its relative efficiency with that of the Neyman-Pearson optimal test which assumes oracle knowledge of $F$, $G$ and $\theta$. This analysis reveals that Shiraishi’s test is most powerful for balanced calibration and test sample sizes, which provides a potentially useful practical guideline. Moreover, it shows that the asymptotic relative efficiency (ARE), or Pitman efficiency, of Shiraishi’s LMPR test relative to the LRT optimal test equals $1-\lambda$ if $n/N\rightarrow\lambda\in(0,1)$.
2.3.3 Asymptotic null distribution of Shiraishi’s test without independence
The critical value $c^{\text{g}}_{\alpha}(m,n)$ in (5) may be obtained from the permutation distribution or from a large-sample approximation. The following theorem extends the asymptotic approximation of the null distribution of $T^{\text{g}}$ given by Shiraishi (1985), which assumed i.i.d. scores, to the weaker assumption of exchangeable scores. This extension is useful because it allows Shiraishi’s test to be applied in combination with the ACODE methodology presented in this paper, which involves a key step that breaks independence while maintaining exchangeability. The proof, in Appendix A5, follows from a general formulation of the finite-population central limit theorem (Hájek, 1960; Li and Ding, 2017).
@@ 2.3.3 Asymptotic null distribution of Shiraishi’s test without independence
As $N\rightarrow\infty$, suppose that
$\frac{\max_{1\leq r\leq N}(\mathbb{E}[g(U_{N}^{(r)})]-\mu_{N})^{2}}{\min(n,m)\sigma^{2}_{N}}\rightarrow 0,$ (6)
where
$\mu_{N}=\frac{1}{N}\sum_{r\in[N]}\mathbb{E}[g(U_{N}^{(r)})],\qquad\sigma^{2}_{N}=\frac{1}{N-1}\sum_{r\in[N]}(\mathbb{E}[g(U_{N}^{(r)})]-\mu_{N})^{2}.$
Then, under the null hypothesis $H^{*}_{0}:\mathrm{the\,\,vector\,\,}(X_{1},\ldots,X_{m},Y_{1},\ldots,Y_{n})\,\,\mathrm{is\,\,exchangeable},$ the standardized Shiraishi statistic converges in distribution to a standard normal,
$\frac{T^{\mathrm{g}}-n\mu_{N}}{\sqrt{\frac{mn}{N}\sigma^{2}_{N}}}\xrightarrow{d}N(0,1).$ (7)
Note that as $N\rightarrow\infty$, condition (6) implies $m\rightarrow\infty$ and $n\rightarrow\infty$ (Li and Ding, 2017). Moreover, the asymptotic approximation of the null distribution of $T^{\text{g}}$ given in Shiraishi (1985, Theorem 1) requires that $g$ be bounded, which in turn implies (6).
As a consequence of Theorem 1, when considering any $S\subseteq[n]$ and denoting by $s$ its cardinality, under the null hypothesis $H_{S}:\cap_{j\in S}H_{j}$, the test statistic
$T^{\text{g}}_{S}$ $=\sum_{j=m+1}^{m+s}\mathbb{E}[g(U_{m+s}^{(R_{j}^{S})})],$ (8)
where $R^{S}_{1},\ldots,R^{S}_{m+s}$ denote the ranks of $(X_{1},\ldots,X_{m},(Y_{j})_{j\in S})$ and $U_{m+s}^{(R_{j}^{S})}$ is the $R^{S}_{j}$-th order statistic in a sample of Uniform random variables of size $m+s$, is asymptotically normal with mean
$s\mu_{m+s}=\frac{s}{m+s}\sum_{r\in[m+s]}\mathbb{E}[g(U^{(r)}_{m+s})$]
and variance
$\frac{ms}{m+s}\sigma^{2}_{m+s}=\frac{ms}{(m+s)(m+s-1)}\sum_{r\in[m+s]}\left(\mathbb{E}[g(U^{(r)}_{m+s})]-\mu_{m+s}\right)^{2}.$
2.3.4 Shirashi’s Adaptive Rank Test
The oracle procedure studied in Theorem 1 can be translated into a practical test by replacing the unknown density $g$ with a suitable empirical estimate $\hat{g}$. In Appendix A2, we establish finite-sample validity of the resulting adaptive rank test (Proposition A3) and derive an asymptotic approximation to its null distribution (Theorem A2).
The following theorem, proved in Appendix A5, implies that Shiraishi’s adaptive rank test is asymptotically equivalent to LMPR test in terms of local asymptotic power. Specifically, we first show that the difference between the oracle and adaptive rank statistics is negligible under the null hypothesis. By contiguity, this negligibility extends to local alternatives, and it follows that the adaptive rank test attains the same local asymptotic power function as the LMPR test with oracle knowledge of $g$. To simplify notation, we replace $\mathbb{E}[g(U^{(r)}_{N})$] by the approximation $g\!\left(\frac{r}{N+1}\right)$. The difference $\mathbb{E}[g(U^{(r)}_{N})]-g\!\left(\frac{r}{N+1}\right)$ is asymptotically negligible and does not affect the asymptotic results.
@@ 2.3.4 Shirashi’s Adaptive Rank Test
Let $X_{1},\ldots,X_{m}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}F$ and $Y_{1},\ldots,Y_{n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}(1-\theta)F+\theta G(F)$, where $F$ is continuous and $G$ admits a continuous and bounded density $g$ on $[0,1]$. Let $N=m+n$ and assume that $n/N\to\lambda\in(0,1)$ as $N\to\infty$.
Let $\hat{g}_{N}$ be an estimate of $g$ based on the unordered pooled sample $\{W_{1},\ldots,W_{N}\}$. If
$\frac{1}{N}\sum_{r=1}^{N}\Bigl[\hat{g}_{N}\!\left(\frac{r}{N+1}\right)-g\!\left(\frac{r}{N+1}\right)\Bigr]^{2}\;\xrightarrow{\mathbb{P}}\;0,$ (9)
then, under the null hypothesis $H_{0}:\theta=0$,
$\frac{T^{\hat{\text{g}}}-n\hat{\mu}_{N}}{\sqrt{N}}-\frac{T^{\text{g}}-n\mu_{N}}{\sqrt{N}}\;\xrightarrow{\mathbb{P}_{0}}0,$
where $T^{\hat{\text{g}}}=\sum_{j=m+1}^{N}\hat{g}_{N}\!\left(\frac{R_{j}}{N+1}\right)$ and $\hat{\mu}_{N}=N^{-1}\sum_{r\in[N]}\hat{g}_{N}\!\left(\frac{r}{N+1}\right)$.
Corollary 1.
As a consequence of Theorem 2, the Shiraishi’s adaptive rank statistic attains the same local asymptotic power function as the oracle LMPR statistic.
In Appendix A2.5, we show that our data-driven estimate of $\hat{g}$ can be implemented using the approach of Patra and Sen (2016), in which case condition (9) is satisfied. Thus, we can leverage an adaptive rank test with the same local asymptotic power as the oracle LMPR test. Moreover, this test can be practically applied within our closed testing methodology described below thanks to the closed-testing shortcut described in Appendix A3.4.
The desirable theoretical properties of this adaptive test are corroborated by the empirical results presented in Section 3, where we show that in some scenarios of interest this adaptive test can achieve much higher power compared to other local testing methods, such as Simes’, Fisher’s, and the WMW test. In any case, the true strength of our ACODE method lies in its flexibility, as it does not depend on any single testing procedure but can instead dynamically select the most effective approach based on the data at hand.
@@ 2.4 Estimating the Number of Outliers via Closed Testing
Input : Individual hypotheses $H_{j}$, for all $j\in[n]$; a method $\phi$ for carrying out local tests $\phi_{J}$ of $H_{J}$, for any $J\subseteq[n]$; significance level $\alpha\in(0,1)$.
@@
2For each $J\subseteq[n]$, test $H_{J}$ at level $\alpha$; let $\phi_{J}\in\{0,1\}$ denote the rejection indicator.
3For each $J\subseteq[n]$, adjust the local test by setting $\bar{\phi}_{J}:=\min\{\phi_{K}:K\supseteq J\}$.
4Query $S\subseteq[n]$: compute the lower bound for the number of true discoveries in $S$,
@@
Output : A $(1-\alpha)$ lower bound $d(S)$ for any queried subset $S\subseteq[n]$.
@@
The first step of Algorithm 1 tests all possible intersection hypotheses $H_{J}$ . We refer to these tests as the local tests. In the second step, a multiplicity adjustment is applied to all local tests. The latter results in $H_{J}$ being rejected, setting $\bar{\phi}_{J}=1$ , if and only if $H_{K}$ was locally rejected for all supersets $K$ of $J$. Finally, $d(S)$ is given by (10). This is easiest to explain in the special case of $S=[n]$, when the lower confidence bound for the total number of outliers is given by the difference between the total number of test points, $n$, and the cardinality of the largest subset for which the multiplicity-adjusted local test failed to reject the null hypothesis. For example, if the global null $H_{[n]}$ is rejected but there exists a subset $K\subseteq[n]$ with cardinality |K|=n1|K|=n-1 for which $\phi_{K}=0$, then $d([n])=1$.
@@
If Algorithm 1 is applied using a valid local test $\phi$ satisfying $\mathbb{P}[\phi_{I_{0}}=1\;;\;H_{I_{0}}]\leq\alpha$, then the output lower bounds $d(S)$ satisfy:
@@
The availability of computational shortcuts for Algorithm 1 hinges on the choice of local tests (Goeman et al., 2019, 2021; Tian et al., 2023). See Table A3 in Appendix A3 for a summary of relevant shortcuts and their computational complexity. Appendix A3.1 reviews the shortcut of Goeman et al. (2019), which reduces the cost of Algorithm 1 to $\mathcal{O}(n\log n)$ in the case of Simes local tests. Appendix A3.3 presents a shortcut for an adaptive version of the Simes local test. Appendix A3.4 introduces a shortcut for the Shirashi test described in Section 2.3.2. Appendix A3.5 reviews the strategy of Tian et al. (2023), which provides similar shortcuts for a family of local tests including Fisher’s method and the WMW test.
@@ 2.5 Conformalized Closed Testing
Valid conformity scores can be obtained through various methods. A common approach involves using a one-class classifier trained on an independent set of inliers (Bates et al., 2023). Alternatively, one can employ a binary classifier via positive-unlabeled learning (Marandon et al., 2024). Although the optimal scoring function can, in principle, be approximated through binary classification (Marandon et al., 2024, Lemma 4.3; see also see Appendix A4), in practice there are settings in which one-class classification approaches may perform better (Liang et al., 2024). Another setting in which binary classifiers can be used is when labeled outliers are available (Liang et al., 2024); however, we focus for simplicity on scenarios where all labeled data consist of inliers.
@@
8 Apply $\mathcal{A}$ to evaluate the calibration and test scores $(X_{1},\ldots,X_{m})$ and $(Y_{1},\ldots,Y_{n})$.
9Compute the ranks $R_{1},\ldots,R_{n+m}$ of the pooled sample and the conformal $p$-values.
10Query $S\subseteq[n]$: compute $d(S)$ using Algorithm 1 (ref. Appendix A3) and the closed-testing adjusted $p$-value $\hat{p}(S)$ for the local test $\bar{\phi}_{S}$ of $H_{S}$.
@@
Output : A $(1-\alpha)$ lower bound $d(S)$ for the number of outliers in $S$, and a $p$-value $\hat{p}(S)$ for testing $H_{S}$, for any queried subset $S\subseteq[n]$.
@@
Proposition 1 implies Algorithm 2 produces valid inferences as long as the local test $\phi_{I_{0}}$ is valid, and our method is precisely designed to achieve this (ref. Section 2.3). While the lower bound $d(S)$ output by Algorithm 2 is simultaneously valid for all possible subsets $S\subseteq[n]$, in the sense of (11), this does not mean Algorithm 2 needs to explicitly output $d(S)$ for all $S\subseteq[n]$. On the contrary, one would typically apply Algorithm 2 focusing on a particular (but possibly data-driven) choice of $S$, as demonstrated in Section 3. In any case, the components of Algorithm 2 involving the model training and the computation of the conformity scores only need to be applied once, irrespective of $S$.